OpenStudy (anonymous):

find the center and radius of the circle with equation (x + 2)2 + (y + 10)2 = 4

OpenStudy (anonymous):

(x+2)^2 + (y+10)^2 = 4 *****

OpenStudy (anonymous):

(x - h)^2 + (y - k)^2 = r^2 ---> general equation of the circle. where (h, k) is the center and r = Radius (x - 4)^2 + (y + 6)^2 = 16 center at (4 , - 6) r = 4

OpenStudy (anonymous):

how did you get the center?

OpenStudy (anonymous):

The standard equation of a circle with center C(h,k) and radius r is : (x - h)2 + (y - k)2 = r2 (x - 4)2 + (y + 6)2 = 16 this is in the form of (x - h)2 + (y - k)2 = r2 by recognition we can see that: h= -4 and k = 6 this is the center of the circle (-4,6) and the radius is basically the square root of 16 which gives us 4 therefore r (radius of the circle )= 4

OpenStudy (anonymous):

you type realllllly fast:0 and thank you! and if you dont mind can you check something for me?

OpenStudy (anonymous):